A man is nine times as old as his son. In nine years, he will be three times as old as his son. How old are each of them now?
Answer by TimothyLamb(4379)

---
x = 9y
x + 9 = 3(y + 9)
---
put the system of linear equations into standard form:
---
x = 9y
x + 9 = 3y + 27
---
x - 9y = 0
x - 3y = 18
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = man = 27
y = son = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php